Conditions for Equilibrium. A 600.0 N window washe
Answer #1


weight of the window washer, w1=600 N at x=2m


weight of the scaffold, w2=450 N


length of the scaffold, l=3m


by balancing force,

T1+T2=w1+W2


T1+T2=600+450


T1+T2=1050 ----(1)


and


net torque w r to left end


T2*l=w2*l/2+w1*(l-x)


T2*3=450*3/2+600*(3-2)

T2=425 N --------------->


from equation (1),


T1+425=1050


T1=625 N ---------------->